OpenStudy (anonymous):

x^2-2x-15 less than or equal to 0

hero (hero):

\[x^2-5x+3x-15 \le 0\]\[x(x-5)-3(x-5) \le 0\]

hero (hero):

(x-5)(x-3) le 0

OpenStudy (anonymous):

\[x ^{2}-2x-15=(x-5)(x+3)\le0\]

OpenStudy (anonymous):

Only when the graphs falls in this interval: \[-3\le x \le 5\]